### v2

@@ 2.2 Chemical language modeling
Language models have been adapted and used for a range of chemical tasks [caldasramos_review_2025]. These models typically operate on linearized molecular strings, such as SMILES [weininger1988smiles], SELFIES [krenn2020self], or IUPAC names. Masked pre-training approaches (ChemBERTa [chithrananda2020chemberta], MolBERT [Fabian2020MolecularRL]) learn molecular fingerprints that are useful for QSAR, whereas the Molecular Transformer family targets forward and retrosynthesis prediction [schwaller2019molecular, schwaller2020predicting, irwin2022chemformer]. More recently, LLMs have been adapted and applied for tackling chemical tasks [Frey2023NeuralSO, zhang2024chemllmchemicallargelanguage, Jablonka2024LeveragingLL, Xie2023DARWINSD]; extending general LLMs for molecule generation, property prediction, and Q&A. Other works have adapted LLMs for use as chemistry agents, integrating robotic labs and other tools [Bran2023AugmentingLL, boiko2023autonomous], hypothesis generation [moosechem, airesearcher, kosmos], and more recently, workflows have been designed for molecular design and synthesis planning [wang2024molleo, bran2025chemical, synthelite, larcretro].
@@ 2.3 Mid-stage domain adaptation
Continuing pre-training on an in-domain corpus—often called domain-adaptive pre-training (DAPT) or continued pre-training (CPT)—has become the dominant recipe for turning a general LLM into a domain specialist. Early successes such as BioMegatron for biomedicine [shin2020biomegatron], Meditron [meditron], Legal-BERT [chalkidis2020legal], and Code-Llama for programming [roziere2023code] demonstrated sizable gains with only a few billion extra tokens. A recent wave of work scales the idea to scientific domains: (1) AdaptLLM [cheng2023adapting] shows that a 7B parameters model, after just 10–15B of financial tokens, rivals BloombergGPT (50B parameters) on in-domain QA; (2) Tag-LLM [shen2024tag] and Efficient-CPT [xie2024efficientcpt] report similar jumps while using parameter-efficient adapters; (3) SciLitLLM [li2024scilitllm] uses a 12.7B token corpus of textbooks and full-text papers and beats much larger baselines on scientific-literature understanding; (4) domain-specific studies in materials science [lu2025finetuning], radiation oncology [holmes2023radiation], Japanese finance [hirano2024finance], and cybersecurity [bayer2024cysecbert] confirm that CPT injects latent domain knowledge that survives further instruction tuning. However, two caveats emerge: CPT can erode zero-shot prompting ability if done naively [cheng2023adapting], and very small models (<< 2B parameters) often fail to develop new capabilities even after extensive CPT [lu2025finetuning, hsieh2025continual]. Crucially, none of these works evaluate whether the adapted model becomes latently solvable for multi-step reasoning tasks that reinforcement learning could later amplify.
The chemical domain remains comparatively under-explored. ChemBERTa-2 [maziarka2023chemberta2] continues a BERT-style encoder on $\sim$1B SMILES tokens and improves fingerprint-style QSAR, while ChemLLM [zhang2024chemllmchemicallargelanguage] , DARWIN [xie2023darwin], and SciDFM [sun2024scidfm], LlaSMol [yu2024llasmol] incorporate reaction patents or literature but still operate in a single-shot, pattern-recognition regime.
@@ 4.1 Datasets
For SFT, we used question-answering (QA) training examples derived from two domains: (1) for organic molecules, we employed SmolInstruct [smolinstruct], utilizing the SMILES $\leftrightarrow$ IUPAC and molecule captioning subsets; (2) for inorganic crystals, we employed MPtrj [deng2023chgnet], utilizing structure to energy/force prediction and initial to relaxed structure mapping subsets, with reasoning traces constructed from intermediate ionic steps to capture the DFT relaxation pathway. Dataset statistics are summarized in Tables 1.Additionally, we incorporated MMLU and Chain-of-Thought (CoT) reasoning traces from DeepSeek-R1, which were preprocessed to maintain coherence with our pretraining data. In this phase, we also expanded the model’s context window from 4,096 to 8,192 tokens to accommodate longer reasoning sequences. The pretrained model underwent SFT for approximately 8 epochs, continuing until the previously observed loss spikes were fully mitigated.
@@ 5 Results
As an initial downstream test of our pipeline’s performance, we use ChemBench [ChemBench] to evaluate the general chemistry knowledge of LLMs; the results are shown in Table 2. This evaluation is a first check to ensure our pipeline consistently improves performance on an important public benchmark for chemistry, before we begin studying the downstream effects of MiST on RL training. The results show an overall +10% improvement in performance bringing it up to 45% which, better than GPT-3.5-Turbo and Llama-3.1-8B-Instruct [mirza2024large], models at least 2x larger than our resulting 3B model. The role of MiST is particularly important in Organic, Inorganic, and General Chemistry, with improvements of up to 6-7% over the Qwen+SFT baseline, and more than 11% over the base (instruction-tuned) model, suggesting benefits of both post-training stages on model’s chemical performance. These results serve already as diagnostic measures of success of a given mid-training methodology, and serve as a basis to select models for the following stage of RLSF, with the goal of enhancing reasoning and problem-solving capabilities.

### v3

@@ 2.2 Chemical language modeling
Language models have been adapted and used for a range of chemical tasks [caldasramos_review_2025]. These models typically operate on linearized molecular strings, such as SMILES [weininger1988smiles], SELFIES [krenn2020self], or IUPAC names. Masked pre-training approaches (ChemBERTa [chithrananda2020chemberta], MolBERT [Fabian2020MolecularRL]) learn molecular fingerprints that are useful for QSAR, whereas the Molecular Transformer family targets forward and retrosynthesis prediction [schwaller2019molecular, schwaller2020predicting, irwin2022chemformer]. More recently, LLMs have been adapted and applied for tackling chemical tasks [Frey2023NeuralSO, zhang2024chemllmchemicallargelanguage, Jablonka2024LeveragingLL, xie2023darwin]; extending general LLMs for molecule generation, property prediction, and Q&A. Other works have adapted LLMs for use as chemistry agents, integrating robotic labs and other tools [Bran2023AugmentingLL, boiko2023autonomous], hypothesis generation [moosechem, airesearcher, kosmos], and more recently, workflows have been designed for molecular design and synthesis planning [wang2024molleo, bran2025chemical, synthelite, larcretro].
@@ 2.3 Mid-stage domain adaptation
Continuing pre-training on an in-domain corpus—often called domain-adaptive pre-training (DAPT) or continued pre-training (CPT)—has become the dominant recipe for turning a general LLM into a domain specialist. Early successes such as BioMegatron for biomedicine [shin2020biomegatron], Meditron [meditron], Legal-BERT [chalkidis2020legal], and Code-Llama for programming [roziere2023code] demonstrated sizable gains with only a few billion extra tokens. A recent wave of work scales the idea to scientific domains: (1) AdaptLLM [cheng2024adaptinglargelanguagemodels] shows that a 7B parameters model, after just 10–15B of financial tokens, rivals BloombergGPT (50B parameters) on in-domain QA; (2) Tag-LLM [shen2024tag] and Efficient-CPT [xie2024efficientcpt] report similar jumps while using parameter-efficient adapters; (3) SciLitLLM [li2024scilitllm] uses a 12.7B token corpus of textbooks and full-text papers and beats much larger baselines on scientific-literature understanding; (4) domain-specific studies in materials science [lu2025finetuning], radiation oncology [holmes2023radiation], Japanese finance [hirano2024finance], and cybersecurity [bayer2024cysecbert] confirm that CPT injects latent domain knowledge that survives further instruction tuning. However, two caveats emerge: CPT can erode zero-shot prompting ability if done naively [cheng2024adaptinglargelanguagemodels], and very small models (<< 2B parameters) often fail to develop new capabilities even after extensive CPT [lu2025finetuning, hsieh2025continual]. Crucially, none of these works evaluate whether the adapted model becomes latently solvable for multi-step reasoning tasks that reinforcement learning could later amplify.
The chemical domain remains comparatively under-explored. ChemBERTa-2 [ahmad2022chemberta2chemicalfoundationmodels] continues a BERT-style encoder on $\sim$1B SMILES tokens and improves fingerprint-style QSAR, while ChemLLM [zhang2024chemllmchemicallargelanguage] , DARWIN [xie2023darwin], and SciDFM [sun2024scidfm], LlaSMol [yu2024llasmol] incorporate reaction patents or literature but still operate in a single-shot, pattern-recognition regime.
@@ 4.1 Datasets
For SFT, we used question-answering (QA) training examples derived from two domains: (1) for organic molecules, we employed SmolInstruct [yu2024llasmol], utilizing the SMILES $\leftrightarrow$ IUPAC and molecule captioning subsets; (2) for inorganic crystals, we employed MPtrj [deng2023chgnet], utilizing structure to energy/force prediction and initial to relaxed structure mapping subsets, with reasoning traces constructed from intermediate ionic steps to capture the DFT relaxation pathway. Dataset statistics are summarized in Tables 1.Additionally, we incorporated MMLU and Chain-of-Thought (CoT) reasoning traces from DeepSeek-R1, which were preprocessed to maintain coherence with our pretraining data. In this phase, we also expanded the model’s context window from 4,096 to 8,192 tokens to accommodate longer reasoning sequences. The pretrained model underwent SFT for approximately 8 epochs, continuing until the previously observed loss spikes were fully mitigated.
@@ 5 Results
As an initial downstream test of our pipeline’s performance, we use ChemBench [mirza2024large] to evaluate the general chemistry knowledge of LLMs; the results are shown in Table 2. This evaluation is a first check to ensure our pipeline consistently improves performance on an important public benchmark for chemistry, before we begin studying the downstream effects of MiST on RL training. The results show an overall +10% improvement in performance bringing it up to 45% which, better than GPT-3.5-Turbo and Llama-3.1-8B-Instruct [mirza2024large], models at least 2x larger than our resulting 3B model. The role of MiST is particularly important in Organic, Inorganic, and General Chemistry, with improvements of up to 6-7% over the Qwen+SFT baseline, and more than 11% over the base (instruction-tuned) model, suggesting benefits of both post-training stages on model’s chemical performance. These results serve already as diagnostic measures of success of a given mid-training methodology, and serve as a basis to select models for the following stage of RLSF, with the goal of enhancing reasoning and problem-solving capabilities.